If the distance between a minima of order 26 and the central maxima is 1215 nm, what is the distance between central maxima and a minima of order 9 in nm? Round to 2 decimals.
Answer #1

We know, distance of nth minima from central maximum, xn = nD/d

Where, = wavelenth, d = slit width, D = distance between screen and slit.

Keeping D,d and constant, xn n

So, for the 26th and 9th minima, we have,

x26/x9 = 26/9 => 1215/x9 = 26/9

=> x9 = (1215*9/26) = 420.58

So, distance for 9th minima = 420.58 nm